Calculate Qsp for calcium fluoride (Ksp = 3.9 ×10–11) when 110.0 mL of a 7.85×10-3 M solution of Ca(NO3)2 is added to 195.0 mL of a 6.75×10-3 M solution of KF.
Answer #1

total volume = (110.0 + 195.0)= 305.0 ml

In the mixture,

[Ca(NO3)2] = [Ca^2+] = 110.0 mL * 7.85×10-3 M / 305.0 ml = 2.83 * 10^-3 M

and

[KF] = [F-] =  195.0 mL * 6.75×10-3 M / 305.0 ml = 4.316 * 10^-3 M

CaF2...............> Ca^2+ + 2 F-

Qsp = [Ca^2+][F-]^2

or

Qsp = 2.83 * 10^-3 * (4.316 * 10^-3)^2

or

Qsp = 5.27 * 10^-8 (answer)